Solve this system of equations by graphing.
-2x+5y=10
y=x+5
A. (-5,0)
B. (5,0)
C. no solution
D. infinitely many solutions

Answer by mananth(16946)

-2 x + 5 y = 10 .............1
-1 x + 1 y = 5 .............2
equation1
when x= 0 y= 2 ( 0 , 2 )
when y = 0 x= -5 ( -5 , 0 )
Equation 2
when x= 0 y= 5 ( 0 , 5 )
when y = 0 x= -5 ( -5 , 0 )

The solution (-5,0)